OpenStudy (anonymous):

y+\frac{ 2 }{ 9 }le \frac{ 7 }{ 8 }

OpenStudy (anonymous):

\[y+\frac{ 2 }{ 9 }\le \frac{ 7 }{ 8 }\]

OpenStudy (anonymous):

subtract \(\frac{2}{9}\) from both sides

OpenStudy (anonymous):

correct